OpenStudy (babyslapmafro):

Please help me get started on the following problem. I am asked to solve the given differential equation by using an appropriate substitution. dy/dx=(y/x)+(x/y)

OpenStudy (babyslapmafro):

\[\frac{ dy }{ dx }=\frac{ y }{ x }+\frac{ x }{ y }\]